there are 10 points ina aplane of which 3 are collinear.how many straight lines wecan drawn ?

Grade:12th Pass

2 Answers

Vaishal Shah
11 years ago
10c2-3c2+1 where c is combination =45-3+1 =43
Aman Bansal
11 years ago

Dear Mangati,

Case 1: We choose two of the four collinear points. Here, only one line is determined because we cannot count 4C2 lines (they are all the same). 

Case 2: We choose two of the six other non-collinear points. Here, it is simply choosing any two points out of 6, 6C2 = 15. 

Case 3: We choose one of the four collinear points, and one of the six remaining points. There are 4*6 = 24 lines determined, no two of which are collinear. 

The number of lines is 1+15+24 = 40. 

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert